Santos Unlimited (SU) was originally unlevered with 4200 shares outstanding. However, after a major financial restructure, SU now has $35000 of debt, with an annual interest expense of 8 percent. The restructuring has reduced the number of shares to 3700. A group of shareholders of SU are not convinced that this move towards adopting financial leverage is a good idea. Their main argument is that there is now some range of EBIT, however low, that will make the shareholders worse off than before. Help understand the situation better by computing the level of earnings before interest and tax (EBIT) that would make shareholders indifferent between being unlevered (i.e. not having any debt) and levered (i.e. having debt). Assume a 31 percent corporate tax rate.

Answer: $ ____
Answer #1

Indifference Level of EBIT or Indifference point-

It is that level of EBIT at which the EPS remains under two different financial plans. At such level of EBIT, it becomes meaningless for the financial manager as to which financial plan or capital structure should be adopted because EPS happens to be the same in each case. In other words, at the indifference level of EBIT, the shareholders would be indifferent whether the funds are raised by one capital mix or another because both will have same level of EPS.

Indifference level of EBIT can be calculated as under:

[EBIT(1-T)]/N1 =  [(EBIT-I)(1-T)]/N2

where,

T = Tax rate

N1 = Number of shares when firm is unlevered

N2 = Number of shares when firm is levered

Interest on debt = 35,000 x 8%

= $2,800

[EBIT(1-0.31)]/4,200 = [(EBIT - 2,800)(1 - 0.31)]/3,700

EBIT/42 = (EBIT - 2,800)/37

37 EBIT = 42 EBIT - 117,600

5 EBIT = 117,600

EBIT = 117,600/5

= $23,520

Hence, at EBIT of $23,520, shareholders would be indifferent between being unlevered and levered.